OpenStudy (anonymous):

Evaluate this exponential expression. (-27)2/3

OpenStudy (jdoe0001):

\(\large{ \bf {\color{blue}{ 27\implies 3\cdot 3\cdot 3\implies 3^3}}\\ \quad \\ (-27)^{\frac{2}{3}}\implies (-3^{\color{blue}{ 3}})^{\frac{2}{3}}\implies -3^{{\color{blue}{ 3}}\cdot \frac{2}{3}}}\)